OpenStudy (anonymous):

Using the function y=x^3−2x^2−3x+4, approximate the slope to the curve at x = 2

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

help please

OpenStudy (loser66):

take derivative, what do you have?

OpenStudy (ranga):

Why are they asking for an approximation? Have you not been taught derivatives yet?

OpenStudy (anonymous):

@Loser66 can you show me what you mean??

OpenStudy (loser66):

to find the slope of the curve, we have to take derivative of the curve and then plug the point in to get the slope of the curve at that point If you don't study derivative yet, you need @ranga help, hehehe because I forgot how to find the slope by approximation way

OpenStudy (ranga):

But cupcakezz would not answer my question if derivative has been taught yet and so I am not able to help.

OpenStudy (anonymous):

@ranga we were, but I don't fully understand it

OpenStudy (anonymous):

can you show me how to get the answer ??

OpenStudy (anonymous):

to approximate, pick another point near x=2 (e.g x = 2.001) and use slope formula to find the slope.

OpenStudy (ranga):

If you have been taught derivatives then you can find the EXACT slope at x = 2. But here they are asking you to approximate the slope at x = 2. Follow sourwing's suggestion above. Approximate slope at x = 2 is: { f(2.001) - f(2) } / 0.001 = ?

OpenStudy (ranga):

f(x) = x^3−2x^2−3x+4 f(2.001) = (2.001)^3 - 2((2.001)^2 - 3(2.001) + 4 = ? f(2) = 2^3 - 2(2)^2 - 3(2) + 4 = ? Then plug it into the formula: Approximate slope at x = 2 is: { f(2.001) - f(2) } / 0.001 = ?

OpenStudy (anonymous):

thanks @ranga